OpenStudy (anonymous):

use comparison theorem/test to determine whether or not intergral [0, infinity] 3/(x+e^x)

OpenStudy (anonymous):

I missed 'converges or diverges' to the question

OpenStudy (anonymous):

\[\int\limits_{0}^{\infty} 3 \div (x +e ^{x} ) dx?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hint from our integral we can have \[\int\limits_{0}^{\infty}1 \div x dx \] which is diverges,

OpenStudy (anonymous):

sorry which diverges,